OpenStudy (anonymous):

Calculus help please? Evaluate this trigonometric integral: integral (sin x/ (1-sin^2 x) ) dx

zepdrix (zepdrix):

\(\Large\bf\sf 1-\sin^2x\quad=\quad \cos^2x\) yes?

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits \frac{\sin x}{\cos^2x}\;dx\quad=\quad \int\limits \frac{1}{\color{royalblue}{\cos}^2\color{royalblue}{x}}\left(\color{orangered}{\sin x\;dx}\right)\]We want to make the substitution: \[\Large\bf\sf \color{royalblue}{u=\cos x}\]Understand how to proceed from there?

OpenStudy (anonymous):

Ok I can see what you did .... do 1/ cos^2x becomes cos x? or does (sin x dx) become cos x?

OpenStudy (anonymous):

@zepdrix ?

zepdrix (zepdrix):

what? I'm not sure what you're asking..

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits\limits \frac{1}{\color{royalblue}{\cos}^2\color{royalblue}{x}}\left(\color{orangered}{\sin x\;dx}\right)\quad=\quad \int\limits\limits \frac{1}{\color{royalblue}{u}^2}\left(\color{orangered}{\sin x\;dx}\right)\]The substitution changes the problem to this, understand? We still need to deal with the sin x dx

OpenStudy (anonymous):

ok I'm sorry now I see

OpenStudy (anonymous):

well sin x dx = - cos x + C right?

zepdrix (zepdrix):

No we don't integrate the sin x dx. We can't integrate something that involves `both` u and x.

zepdrix (zepdrix):

We need to change `everything with x in it` to `something with u in it`. We replaced the cosx with u. Now we need to replace sinx dx with something involving u, BEFORE we can integrate.

zepdrix (zepdrix):

To find a replacement, take the derivative of your substitution,\[\Large\bf\sf \color{royalblue}{u=\cos x},\qquad\qquad\qquad \frac{du}{dx}=?\]

OpenStudy (anonymous):

- sin x?

OpenStudy (anonymous):

or 0 my bad... I think

zepdrix (zepdrix):

Ok good, your first guess looks right. We'll "multiply" the dx to the other side, giving us:\[\Large\bf\sf du=-\color{orangered}{\sin x \;dx}\] See how we `almost` have something to replace with sin x dx in our integral?

OpenStudy (anonymous):

ok so my first guess was right :)

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

We want to "solve for" sin x dx. See how the negative sign is in the way? What can we do about that? :o

OpenStudy (anonymous):

tranfer it to du?

zepdrix (zepdrix):

Multiply the negative to the other side? Ok good.\[\Large\bf\sf \color{orangered}{-du=\sin x \;dx}\]Ok ok ok. See what we can plug in for our sinx dx now?

OpenStudy (anonymous):

hmm I think thats - cos ^-1 ?

zepdrix (zepdrix):

Ok looks good! But `do not` use -1 for powers of trig functions. We reserve that for the inverse function.\[\Large\bf\sf \cos^{-1}x\quad\ne\quad \frac{1}{\cos x}\] You can write it like this if you want,\[\Large\bf\sf (\cos x)^{-1}\] I think your negative goes away when you integrate, yes? You picked up a -1 from the sinxdx substitution. Then you lose it when you integrate, gaining a factor of -1 from the u.

OpenStudy (anonymous):

ok so 1/cos x is my final answer right?

zepdrix (zepdrix):

Mmm ya that sounds right! good job.

OpenStudy (anonymous):

thank you soo much!!